OpenStudy (anonymous):

Square the Binomial (5X-6)^2

OpenStudy (anonymous):

(5x-6) (5x-6) Do you follow me so far?

OpenStudy (anonymous):

You just restated the equation in a different form, (5x-6)^2 is the same as (5x-6)(5x-6)

OpenStudy (anonymous):

this is my first day on here so no Im not following you. how do I simplify that answer?

OpenStudy (anonymous):

thats okay, this equation requires you to FOIL, do you know how to do that?

OpenStudy (anonymous):

no I dont

OpenStudy (anonymous):

To foil, you would multiply the equation against itself. I will do the math and let you see how I do it.

OpenStudy (anonymous):

(5x-6)(5x-6) 5x * 5x = 25x^2 5x * -6 = -35x -6 * 5x = -35x -6 * -6 = 36

OpenStudy (anonymous):

Do you understand how I did that???

OpenStudy (anonymous):

sorry I'm just now responding. how would I simplify (4x-9)^2